# Efficient Data-Sketches and Fine-Tuning for Early Detection of Distributional Drift in Medical Imaging

### v1

@@ Abstract
Distributional drift detection is important in medical applications as it helps ensure the accuracy and reliability of models by identifying changes in the underlying data distribution that could affect diagnostic or treatment decisions. However, current methods have limitations in detecting drift; for example, the inclusion of abnormal datasets can lead to unfair comparisons. This paper presents an accurate and sensitive approach to detect distributional drift in CT-scan medical images by leveraging data-sketching and fine-tuning techniques. We developed a robust baseline library model for real-time anomaly detection, allowing for efficient comparison of incoming images and identification of anomalies. Additionally, we fine-tuned a vision transformer pre-trained model to extract relevant features using breast cancer images as an example, significantly enhancing model accuracy to 99.11%. Combining with data-sketches and fine-tuning, our feature extraction evaluation demonstrated that cosine similarity scores between similar datasets provide greater improvements, from around 50% increased to 100%. Finally, the sensitivity evaluation shows that our solutions are highly sensitive to even 1% salt-and-pepper and speckle noise, and it is not sensitive to lighting noise (e.g., lighting conditions have no impact on data drift). The proposed methods offer a scalable and reliable solution for maintaining the accuracy of diagnostic models in dynamic clinical environments.
@@
Data Drift, Data Sketches, Anomaly Detection, Medical Images
@@ I Introduction
Distributional drift [1], also known as dataset drift, in medical imaging refers to changes in data distribution over time, which can significantly affect the performance of machine learning models used for diagnostic purposes. This drift may result from various factors, including alterations in imaging equipment, differences in imaging protocols, variations in patient demographics, or updates in image preprocessing techniques. Detecting and managing drift is critical in the medical field to ensure that models remain accurate and reliable. Ignoring drift can lead to incorrect diagnoses or suboptimal treatment recommendations, thereby potentially compromising patient care. Therefore, continuous monitoring and adaptation of models are essential to maintain their effectiveness in dynamic clinical environments. Despite significant advancements in data drift detection within medical imaging, current methodologies exhibit several limitations that must be addressed to improve their effectiveness. We discuss the limitations of L1, L2 and L3 methods as follows:
@@
(L1): Analyzing the distribution of image data is particularly challenging due to its high-dimensional characteristics. Unlike numerical data, where statistical properties like mean and variance can be straightforwardly computed and compared, image data necessitates specialized feature extraction and dimensional reduction techniques to perform meaningful distribution analysis.
Figure 1: Med-MNIST Benchmarks. We highlight three abnormal images in the benchmarks.
@@
(L2): Abnormal datasets can introduce bias in drift detection, which can often be mitigated without compromising the overall data quality. However, efficiently identifying and isolating these abnormal datasets is essential to preserve data integrity. Such abnormalities may arise from data entry errors, sensor malfunctions, or even malicious activities. Implementing effective and efficient anomaly detection techniques is critical for pre-processing data before applying drift detection methods. For instance, as shown in Fig. 1, three medical images were identified as unsuitable for model training, thereby impacting the overall data quality.
@@
(L3): Another limitation is the absence of real-time processing solutions for handling abnormal data, especially during the preprocessing stage. This process often relies on hospital experts to manually label and remove invalid data before training, which significantly increases the cost and complexity of model development. When datasets contain substantial noise or inconsistencies, the drift detection process may incorrectly identify or overlook drifts, thereby compromising the reliability of the analysis.
@@
To overcome the identified limitations, we fine-tuned a pre-trained model specifically optimized for feature extraction from the target images. We utilized the breast cancer dataset from MetaStar [2] and applied fine-tuning techniques to the Vision Transformer (ViT)[3] model as an example for pre-train model selection. The selection of ViT was based on its superior accuracy post-fine-tuning, as demonstrated by our experimental results in the evaluation section. The fine-tuning process of ViT involved several crucial steps. Initially, we loaded the model with weights pre-trained on ImageNet [4], a large-scale image dataset, providing a robust foundation for feature extraction. We then replaced the fully connected layers with a custom architecture tailored to our specific task, enabling the model to learn the intricate details of the breast cancer dataset. Additionally, we selectively unfroze the later layers of the model, allowing them to be retrained on our dataset while keeping the earlier layers fixed. This strategy enabled us to fine-tune the model effectively, preserving the general features learned from the larger dataset while adapting it to the unique characteristics of the breast cancer images.
@@
Additionally, our study utilizes advanced data sketching [5, 6] techniques, specifically employing MinHash [7] as the sketching method, to build a robust baseline library for anomaly detection in image pre-processing phase. This baseline model acts as a reference, allowing systematic comparison of incoming images to identify and correct anomalies within the dataset. By leveraging data sketches, the image data is condensed into a smaller, fixed-size structure—the MinHash signature—which can be quickly compared with other sketches to estimate similarity. A low similarity score may indicate abnormal or low-quality images that require further review.
@@
Furthermore, the extracted features from selected images are rigorously compared with the baseline library to detect drift, using Kolmogorov-Smirnov [8] and cosine similarity [9] scores. We evaluate it without data sketches and fine-tuning, comparing it against our proposed methods. These scores are crucial for identifying subtle changes and deviations within the datasets.
@@
The reminder of this paper is organized as follows: Section II provides an overview of the relevant background. Section III details our proposed methods. Section IV presents the experimental results, while Section V discusses related work. Finally, we conclude our research in Section VI.
@@ II-B Causes of Distributional Drift in Image Data
III Proposed Methods
@@
III-A Fine-tuned Pre-trained Model for Feature Extractions
@@ III-A1 Data Preparation
III-A2 Layer Freezing and Fine-Tuning
@@ III-A5 Training Setup
III-B Data Sketches for Real-time Anomaly Detection
@@
III-B1 Feature Extraction and Hashing
@@
III-B2 Constructing the Baseline Library
@@
III-B3 Integration with Real-time Feature Extraction
@@
III-B4 Real-time Anomaly Detection
@@ 2 Hash the feature vector to obtain data sketch $\mathbf{h}_{\text{new}}$;
III-C Similarity Comparisons for Drift Detection
III-C1 Kolmogorov-Smirnov (KS) Statistic
Let $F_{\mathcal{B}}(x)$ and $F_{\text{new}}(x)$ be the empirical CDFs of the baseline library and the new data sketch, respectively. The KS statistic $D$ is defined as:
@@ III-C2 Cosine Similarity
III-D Sensitivity Evaluation via Incremental Noise Introduction
@@
IV Experimental Results
IV-A Experimental Datasets
In our research, we utilize the MedMNIST [13] dataset as benchmark and 9,987 breast cancer images for fine-tuning. The MedMNIST is an extensive repository of standardized biomedical images specifically created to support the development and assessment of machine learning algorithms in medical imaging. MedMNIST mirrors the structure of the MNIST dataset but is significantly larger and specifically adapted for biomedical purposes. It comprises ten distinct 2D datasets and eight 3D datasets. The ten datasets in MedMNIST encompass a range of medical conditions and imaging techniques. For our analysis, we have selected eight common 2D datasets as benchmarks including PathMNIST, BloodMNIST, OrganMNIST, ChestMNIST, DermaMNIST, BreastMNIST, OCTMNIST and PneumoniaMNIST.
@@ IV-A Experimental Datasets
Figure 3: Baseline 1: Kolmogorov-Smirnov Statistic and P-value Trends Across 8 MNIST Datasets.
@@
Figure 4: Baseline 2: Cosine Similarity Score Across 8 Med-MNIST Datasets. (without fine-tuning)
IV-A1 Data Splits
@@
IV-B Evaluation of Model Accuracy
@@
Model Accuracy Fine-Tuned Accuracy
Customized Model 78.7% N/A
ResNet50 N/A 66.9%
ResNet152 N/A 66.7%
VGG16 N/A 92.2%
Vision Transformer(ViT) N/A 99.11%
We evaluated the performance of 5 distinct models—Customized Model, VGG16, and ResNet50, ResNet152 [14] and ViT (patch16)[3]—on breast cancer image classification, as shown in Table I. The primary goal was to assess the effectiveness of fine-tuning when applied to pre-trained models in comparison to a baseline CNN model trained from scratch.
@@
IV-C Evaluation of Drift Detection
Baseline 1 (without data-sketches and fine-tuning). From the experimental results shown in Fig. 3, it is evident that significant data drift starts to emerge around the sixth month. This drift is reflected by the noticeable increase in the Kolmogorov-Smirnov (KS) statistic across most of the MNIST dataset variants, such as PathMNIST, BloodMNIST, and others. This suggests that changes in the underlying data distribution begin at this point, which could negatively impact model performance.
Although there is some level of similarity between datasets over time, this similarity is unstable. The fluctuations in the KS statistic across different time points indicate inconsistency in the relationship between the datasets. For certain months, the data distributions may remain relatively stable, while in others, there are substantial shifts. This instability makes it challenging to draw reliable conclusions from the observed similarities, especially in a sensitive domain like medical imaging. Moreover, the absolute values of the KS statistic show that the overall similarity scores are relatively low. Without fine-tuning, the feature extraction process struggles to capture the deeper similarities between datasets, resulting in poor comparison scores. This indicates that the raw features lack sufficient discriminatory power when comparing different datasets, especially when significant data drift occurs. The inability of the non-fine-tuned model to adequately perceive feature differences highlights the necessity of fine-tuning to better adapt to the specific characteristics of the datasets and improve the accuracy and stability of the feature comparisons.
Figure 5: Baseline 3: Cosine similarity of fine-tuned model without data-sketches. (fine-tuned without data-sketches and noises)
@@ IV-C Evaluation of Drift Detection
Our Performance (fine-tuned with data-sketches). In Fig. 6, the cosine similarity scores using BreastMNIST with varying levels of noise demonstrate that the feature extraction process is of high quality. For the first three monthly datasets, the similarity scores highly stable (100% similarity with no fluctuations) across different datasets, indicating no significant drift and consistent feature representation. This stability reflects our solution’s robustness in handling abnormal and bad quality datasets.
IV-D Sensitivity Evaluation
@@
Figure 6: Our Performance in Breast Pre-train Model with Data-sketches: Cosine Similarity Score Across 8 Random Noises.
@@
In the background, Section II-B, we introduced several factors that can cause dataset drift. Here, we introduce four types of noise—Gaussian Noise, Salt-and-Pepper Noise, Speckle Noise, and Poisson Noise [15]—each corresponding to one of these drift-inducing factors. These noises are used to test the sensitivity of our method under different conditions that may simulate real-world drift scenarios
IV-D1 Gaussian Noise

### v2

@@ Abstract
Distributional drift detection is important in medical applications as it helps ensure the accuracy and reliability of models by identifying changes in the underlying data distribution that could affect the prediction results of machine learning models. However, current methods have limitations in detecting drift, for example, the inclusion of abnormal datasets can lead to unfair comparisons. This paper presents an accurate and sensitive approach to detect distributional drift in CT-scan medical images by leveraging data-sketching and fine-tuning techniques. We developed a robust baseline library model for real-time anomaly detection, allowing for efficient comparison of incoming images and identification of anomalies. Additionally, we fine-tuned a pre-trained Vision Transformer model to extract relevant features, using mammography as a case study, significantly enhancing model accuracy to 99.11%. Combining with data-sketches and fine-tuning, our feature extraction evaluation demonstrated that cosine similarity scores between similar datasets provide greater improvements, from around 50% increased to 99.1%. Finally, the sensitivity evaluation shows that our solutions are highly sensitive to even 1% salt-and-pepper and speckle noise, and it is not sensitive to lighting noise (e.g., lighting conditions have no impact on data drift). The proposed methods offer a scalable and reliable solution for maintaining the accuracy of diagnostic models in dynamic clinical environments.
@@
Data Drift, Image Quality, Anomaly Detection, Medical Images
@@ I Introduction
Distributional drift [1], also known as dataset drift, in medical imaging refers to changes in data distribution over time, which can significantly affect the accuracy of machine learning models used for diagnostic purposes. This drift may result from various factors, including alterations in imaging equipment, differences in imaging protocols, variations in patient demographics, or updates in image preprocessing techniques. Detecting and managing drift is critical in the medical field to ensure that models remain accurate and reliable. Ignoring drift can lead to incorrect diagnoses or suboptimal treatment recommendations, thereby potentially compromising patient care. Therefore, continuous monitoring and adaptation of models are essential to maintain their effectiveness in dynamic clinical environments. Despite significant advancements in data drift detection within medical imaging, current methodologies exhibit several limitations that must be addressed to improve their effectiveness. We discuss the limitations of L1, L2 and L3 methods as follows:
@@
(L1): Analyzing the distribution of image data is particularly challenging due to its high-dimensional characteristics. Unlike numerical data, where statistical properties like mean and variance can be straightforwardly computed and compared, image data needs specialized feature extraction and dimensional reduction techniques to perform meaningful distribution analysis.
Figure 1: Med-MNIST Benchmarks. We highlight three unqualified images in the benchmarks.
@@
(L2): Abnormal datasets can introduce bias in drift detection, which can often be mitigated without compromising the overall data quality. However, efficiently identifying and isolating these abnormal datasets is essential to preserve data integrity. Such abnormalities may arise from data entry errors, sensor malfunctions, or even malicious activities. Implementing effective and efficient anomaly detection techniques is critical for pre-processing data before applying drift detection methods. For instance, as shown in Fig. 1, three medical CT images were identified as unqualified for model training.
@@
(L3): The absence of real-time processing solutions to handle abnormal data, especially during the pre-processing stage. This process often relies on hospital experts to manually label and remove invalid data before training, which significantly increases the cost and complexity of model development. A solution that enables fast detection is needed.
@@
To overcome the identified limitations, we fine-tuned a pre-trained model specifically optimized for feature extraction from the target images. We fine-tuned Vision Transformer (ViT)[2] model with 9,987 mammography images as a case study. The selection of ViT was based on its superior accuracy post-fine-tuning, as demonstrated by our experimental results in the evaluation section. The fine-tuning process of ViT involved several crucial steps. Initially, we loaded the model with weights pre-trained on ImageNet [3], a large-scale image dataset, providing a robust foundation for feature extraction. We then replaced the fully connected layers with a custom architecture tailored to our specific task, enabling the model to learn the intricate details of the breast cancer dataset. Additionally, we selectively unfroze the later layers of the model, allowing them to be retrained on our dataset while keeping the earlier layers fixed. This strategy enabled us to fine-tune the model effectively, preserving the general features learned from the larger dataset while adapting it to the unique characteristics of the breast cancer images.
@@
Additionally, our study designed an advanced data sketching [4, 5] solution, specifically employing MinHash [6] as the sketching method, to build a robust baseline library for anomaly detection in image pre-processing phase. This baseline model acts as a reference, allowing systematic comparison of incoming images to identify and correct anomalies within the dataset. By leveraging data sketches, the image data is condensed into a smaller, fixed-size structure, the MinHash signature, which can be quickly compared with other sketches to estimate similarity. A low similarity score may indicate abnormal or low-quality images that require further review.
@@
Finally, after excluding all the low score images, all the extracted features from selected images are rigorously compared with the baseline library to detect drift, using both Kolmogorov-Smirnov [7] and cosine similarity [8] algorithms. Our solutions compared with 3 baseline models. The results show that our solution is efficient in identifying subtle changes and deviations within the medical images.
@@
The reminder of this paper is organized as follows: Section II provides an overview of the relevant background. Section IV details our proposed methods. Section V presents the experimental results, while Section III discusses related work. Finally, we conclude our research in Section VI.
@@ II-B Causes of Distributional Drift in Image Data
III Related Work
Existing research on distributional drift detection has predominantly focused on textual data, leveraging changes in word embeddings [9] or similar features to monitor drift [10, 11, 12]. While these methods have demonstrated effectiveness in text-based applications, they face significant challenges when applied to high-dimensional medical imaging data. Unlike text, subtle variations in pixel intensities or textures in images can have substantial clinical implications, demanding more sensitive and robust approaches.
Several studies have explored transfer learning to adapt pre-trained models to new datasets [13, 14, 15, 16]. Transfer learning effectively reduces the need for large training datasets, but it is prone to errors from noisy or outlier data. The reliance on models that are highly sensitive to such imperfections limits their capacity to detect minor distributional changes, which is crucial in medical applications. Furthermore, these approaches often overlook real-time requirements, making them unsuitable for dynamic clinical environments where prompt drift detection is critical.
Drift detection in medical imaging remains underexplored compared to text data. Existing methods typically involve visual inspection or traditional image processing techniques [17, 18, 19], which are often not scalable for large datasets or capable of handling the complexity of modern imaging data. For instance, approaches relying solely on statistical tests or basic similarity measures fail to capture deeper relationships within high-dimensional image datasets, leading to inconsistent and unreliable drift detection results.
Our proposed method addresses these limitations by integrating fine-tuning of a ViT model with advanced data sketching techniques. This combination enhances the robustness and sensitivity of drift detection, even in the presence of minimal noise. Unlike methods that rely heavily on statistical metrics alone, our approach leverages the strengths of data sketches to efficiently summarize high-dimensional data and employs cosine similarity to identify subtle distributional shifts. Additionally, by incorporating real-time anomaly detection mechanisms, our method is well-suited for dynamic clinical settings, offering a scalable and reliable solution for maintaining diagnostic model performance.
In contrast to prior approaches, which often struggle with noisy or imbalanced datasets, our method demonstrates resilience by filtering low-quality data and optimizing feature extraction. By achieving stable similarity comparisons and high sensitivity to minor anomalies, our solution ensures more accurate drift detection, addressing key gaps in existing methodologies.
IV Proposed Methods
@@
IV-A Fine-tuned Pre-trained Model for Feature Extractions
@@ IV-A1 Data Preparation
IV-A2 Layer Freezing and Fine-Tuning
@@ IV-A5 Training Setup
IV-B Data Sketches for Real-time Anomaly Detection
@@
IV-B1 Feature Extraction and Hashing
@@
IV-B2 Constructing the Baseline Library
@@
IV-B3 Integration with Real-time Feature Extraction
@@
IV-B4 Real-time Anomaly Detection
@@ 2 Hash the feature vector to obtain data sketch $\mathbf{h}_{\text{new}}$;
IV-C Similarity Comparisons for Drift Detection
IV-C1 Kolmogorov-Smirnov Statistic
The Kolmogorov-Smirnov (KS) test is a non-parametric statistical method used to compare the cumulative distribution functions (CDFs) of two datasets or a dataset and a theoretical distribution. It evaluates the maximum distance between the distributions (D-statistic) and determines whether the observed difference is statistically significant. The p-value associated with the KS test indicates the likelihood of observing the difference under the null hypothesis, which assumes the datasets come from the same distribution. A low p-value (e.g., $<$ 0.05) suggests a significant difference, while a high p-value indicates no evidence to reject the null hypothesis, implying similarity between distributions.
Formal definition: Let $F_{\mathcal{B}}(x)$ and $F_{\text{new}}(x)$ be the empirical CDFs of the baseline library and the new data sketch, respectively. The KS statistic $D$ is defined as:
@@ IV-C2 Cosine Similarity
IV-D Sensitivity Evaluation via Incremental Noise Introduction
@@
V Experimental Results
V-A Experimental Datasets
In our research, we utilize the MedMNIST [23] dataset as benchmark and 9,987 mammography images for fine-tuning. The MedMNIST is an extensive repository of standardized biomedical images specifically created to support the development and assessment of machine learning algorithms in medical imaging. MedMNIST mirrors the structure of the MNIST dataset but is significantly larger and specifically adapted for biomedical purposes. It comprises ten distinct 2D datasets and eight 3D datasets. The ten datasets in MedMNIST encompass a range of medical conditions and imaging techniques. For our analysis, we have selected eight common 2D datasets as benchmarks including PathMNIST, BloodMNIST, OrganMNIST, ChestMNIST, DermaMNIST, BreastMNIST, OCTMNIST and PneumoniaMNIST.
@@ V-A Experimental Datasets
Figure 3: Baseline 1 (without fine-tuning): KS Statistic and P-value (IV-C1) Trends Across 8 Med-MNIST Datasets. P-value shows that all the datasets are $<$ 0.05, meaning the low similarities in data distribution. KS statistics trends for the BreastMNIST dataset highlights a critical challenge in identifying data drift within highly variable datasets. Unlike other MNIST variants, the trends in BreastMNIST exhibit significant fluctuations across time, which complicates the interpretation of when drift occurs. The instability in KS statistics suggests that the dataset may be influenced by factors such as inconsistencies in data acquisition processes or intrinsic variability in the imaging data.
@@
Figure 4: Baseline 2 (without fine-tuning): Cosine Similarity Score (IV-C2) Across 8 Med-MNIST Datasets. The analysis of cosine similarity scores across the 8 Med-MNIST datasets reveals a fundamental limitation of feature extraction without fine-tuning. The similarity scores remain consistently low, hovering around 50%, which indicates that the extracted features fail to capture meaningful relationships between datasets. This low level of similarity suggests that the pre-trained model, when used without fine-tuning, is unable to adapt to the specific characteristics of medical imaging data.
V-A1 Data Splits
@@
V-B Evaluation of Model Accuracy
@@
Model Fine-Tuned Accuracy
Customized Model 78.7%
ResNet50 66.9%
ResNet152 66.7%
VGG16 92.2%
Vision Transformer(ViT) 99.11%
We evaluated the performance of 5 distinct models, Customized Model, VGG16, and ResNet50, ResNet152 [24] and ViT (patch16)[2], on breast cancer image classification, as shown in Table I. The primary goal was to assess the effectiveness of fine-tuning when applied to pre-trained models in comparison to a baseline CNN model trained from scratch.
@@
V-C Real-time Detection with New Images
Figure 5: Real-time image similarity comparison
Fig. 5 shows individual images compared with the baseline sketch model. Each image is assigned a similarity score, where a lower score may suggest dissimilar images that require further inspection. Therefore, such images should be highlighted, and the doctor must manually define an appropriate threshold to identify and review them.
V-D Evaluation of Drift Detection
Baseline 1 (without data-sketches and fine-tuning) Fig. 3 reveals significant trends in the KS statistics and p-values across multiple datasets, providing insights into the onset and progression of data drift. From Fig. 3, it is evident that the KS statistic for BreastMNIST increases steadily over time, starting from approximately 0.003 in the third month and peaking at 0.007 by the sixth month. This upward trend strongly suggests a substantial change in the underlying data distribution.
For PathMNIST and BloodMNIST, a similar rise in the KS statistic is observed, with values exceeding 0.005 around the fourth month. These consistent increases indicate the emergence of distributional drift across these datasets. Importantly, the corresponding p-values for all datasets remain consistently below 0.05, and in some cases, such as BreastMNIST, they fall below 0.01, indicating that the observed differences in data distributions are statistically significant.
The variability in the KS statistic trends across datasets highlights the complex nature of data drift in medical imaging. For BreastMNIST, the KS statistic shows substantial fluctuations month-to-month, which could be attributed to factors such as variations in imaging protocols, patient demographics, or inherent noise in the data collection process. In contrast, datasets like PathMNIST exhibit a more gradual and consistent increase in KS values, reflecting a smoother drift pattern.
The consistently low p-values ($<$ 0.05) further confirm that the detected distributional differences are unlikely to be due to random chance. For example, in the sixth month, the p-value for BloodMNIST drops to approximately 0.008, reinforcing the significance of the observed drift. These findings suggest that the datasets are diverging from their initial distributions in ways that are both measurable and meaningful.
The analysis of KS statistics and p-values from Fig. 3 underscores the importance of robust drift detection mechanisms. While the KS test effectively highlights significant distributional changes, the observed fluctuations and consistently low p-values point to the limitations of raw feature extraction methods. These results emphasize the need for fine-tuning and additional techniques to better adapt to the specific characteristics of each dataset and ensure reliable drift detection.
Figure 6: Baseline 3 (fine-tuned without data-sketches and noises): Cosine similarity of fine-tuned model without data-sketches. We fine-tuned the BreastMNIST ViT model; however, without applying data sketches to remove noise, the similarity scores still fluctuate, making it difficult to reliably detect data drift.
@@ V-D Evaluation of Drift Detection
Figure 7: Our Performance in Breast Pre-train Model with Data-sketches: Cosine Similarity Score Across 8 Random Noises. The evaluation shows that our method maintains high cosine similarity scores (99%) under low noise levels, demonstrating stability and robustness. As noise increases, the scores decline, but the method effectively detects shifts, showcasing its sensitivity to distributional changes.
@@
Our Performance (fine-tuned with data-sketches). In Fig. 7, the cosine similarity scores using BreastMNIST with varying levels of noise demonstrate that the feature extraction process is of high quality. For the first three monthly datasets, the similarity scores highly stable (around 99% similarity with no fluctuations) across different datasets, indicating no significant drift and consistent feature representation. This stability reflects our solution’s robustness in handling abnormal and bad quality datasets.
V-E Sensitivity Evaluation
In the background, Section II-B, we introduced several factors that can cause dataset drift. Here, we introduce four types of noise: Gaussian Noise, Salt-and-Pepper Noise, Speckle Noise, and Poisson Noise [25], each corresponding to one of these drift-inducing factors. These noises are used to test the sensitivity of our method under different conditions that may simulate real-world drift scenarios
V-E1 Gaussian Noise